OpenStudy (minime):

2x+5y= 2 3x-2y= -16 please solve step by step using elimanation

OpenStudy (minime):

HELP

OpenStudy (anonymous):

multiply top by 3 and bottom by -2 so 6x+15y=6 -6x+4y=32 so 19y=38 y=2 and x=-4